OpenStudy (pandabear500):

What is the standard form of equation of a circle with center at (-2, -3) and passing through the point (-2, 0)? A. (x + 2)2 + (y + 3)2 = 9 B. (x − 2)2 + (y − 3)2 = 16 C. (x − 2)2 + (y − 3)2 = 4 D. (x + 2)2 + (y + 3)2 = 16 E. (x − 2)2 + (y + 3)2 = 9

OpenStudy (welshfella):

general form is (x - a)^2 + (y - b)^2 = r^2 where (a,b) is the center and r = radius.

OpenStudy (welshfella):

so if you replace a with -2 and y with -3 and simplify you'll be able to see which ones of the choices it could be.

OpenStudy (welshfella):

(x - (-2))^2 + (y - (-3))^2 = r^2 what does -(-2) and -(-3) simplify to?

OpenStudy (welshfella):

- 2 negatives together like this make a plus

OpenStudy (welshfella):

so ( x --2) = (x + 2)

OpenStudy (welshfella):

do you follow what I mean?

OpenStudy (welshfella):

dont worry about the r^2 at the moment can you work out what the left part of the equation will be?

OpenStudy (welshfella):

r u there?

OpenStudy (pandabear500):

You kinda helped but I don't really think you understand it. But I understand it alil better now so thank you @welshfella

OpenStudy (welshfella):

Oh i understand it alright. from the above reasoning you'll find its 1 of 2 choices. Then to find the correct choice you can plug in the point (-2 ,0) - that is x = -2 and y = 0 and see which one fits

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

can you see what left side of the equation will be it will start with (x + 2)^2